OpenStudy (anonymous):

what is the solution of 1/2x - 8 = 3/2 x + 2 please show work, thank yall so much.

OpenStudy (tcp2315):

0.5x-8 = 1.5x +2 0.5x-8 (-0.5x) =1.5x+2 (-0.5x) -8 (-2) = 1x +2 (-2) -10= x